# HL-IK: A Lightweight Implementation of Human-Like Inverse Kinematics in Humanoid Arms

### v2

@@ IV-C Teleoperation Experiments on a Physical Robot
We tested a robot model in the physical world that is completely different from the one used in simulation—a wheeled-arm robot’s upper limb. We use the same weight coefficients as those used in the simulation. We applied the HL-IK method to the robot’s left arm and the original IK method to the right arm, observing the differences in arm configurations during the teleoperation task. We use a Meta VR headset as the sensing input for the EE target, and no other sensing is employed. As shown in Fig. 9, HL-IK enabled the robot to track the end effector and, without direct sensing, infer and align with the human elbow position. As a result, the robot’s overall arm configuration closely matched the human’s, whereas the baseline IK achieved only end effector tracking and failed to reproduce configuration level similarity. In particular, this difference becomes increasingly apparent with higher elbow joint heights.
V Conclusions
In this paper, we introduced HL-IK, a lightweight IK framework that preserves EE tracking while producing human-like arm configurations without full-body sensing. The key idea is a learned elbow prior: a FiSTA network predicts the next-step elbow pose from a short motion history and the desired hand target, and this prediction is injected as a small residual cost alongside EE tracking and smoothness in a standard Levenberg–Marquardt IK stack. This design integrates seamlessly with generic kinematics libraries, adds negligible overhead (best trade-off at a 5-frame history), and significantly improves anthropomorphism. On a 183k-step test, HL-IK reduces arm key-point and directional errors by 30.6% and 35.4% on average across all steps—and by 42.2% and 47.4% on the most challenging trajectories—while maintaining acceptable EE accuracy. Teleoperation experiments on a physical robot likewise clearly demonstrate the effectiveness of our method in improving arm-configuration similarity in IK. Looking ahead, we will explore richer upper-limb priors, extend to whole-body control, enable online adaptation and personalization, and unify collision and task constraints within a single optimization framework to further improve generalization and robustness in real deployments.

### v3

@@ IV-C Teleoperation Experiments on a Physical Robot
We tested HL-IK on a physical robot that differs substantially from the simulated platform-the upper limbs of a wheeled mobile manipulator. We deployed HL-IK on the left arm and a Jacobian-based numerical IK on the right arm, and qualitatively compared their arm configurations during teleoperation. Our hardware study primarily aims to demonstrate end-to-end teleoperation feasibility and user experience; therefore, we benchmark against a Jacobian-based IK that is widely used in practical teleoperation for robots, rather than the swivel-angle–based IK evaluated in simulation. We use a Meta VR headset to provide the EE target, and no additional sensing is used. As shown in Fig. 9, HL-IK enabled the robot to track the end effector and, without direct sensing, infer and align with the human elbow position. As a result, the robot’s overall arm configuration closely matched the human’s, whereas the baseline IK achieved only end effector tracking and failed to reproduce configuration level similarity. In particular, this difference becomes increasingly apparent with higher elbow joint heights.
V Conclusions and Limitations
In this paper, we introduced HL-IK, a lightweight IK framework that preserves EE tracking while producing human-like arm configurations without full-body sensing. The key idea is a learned elbow prior: a FiSTA network predicts the next-step elbow pose from a short motion history and the desired hand target, and this prediction is injected as a small residual cost alongside EE tracking and smoothness in a standard Levenberg–Marquardt IK stack. This design integrates seamlessly with generic kinematics libraries, adds negligible overhead (best trade-off at a 5-frame history), and significantly improves anthropomorphism. On a 183k-step test, HL-IK reduces arm key-point and directional errors by 30.6% and 35.4% on average across all steps—and by 42.2% and 47.4% on the most challenging trajectories—while maintaining acceptable EE accuracy. Teleoperation experiments on a physical robot likewise clearly demonstrate the effectiveness of our method in improving arm-configuration similarity in IK.
Despite these encouraging results, our current formulation primarily leverages elbow alignment. While effective in improving overall arm appearance, human-likeness is also influenced by other factors such as shoulder shrugging, torso compensation and habitual wrist posture. Consequently, using only an elbow residual may not consistently match human subjective preferences across all tasks, particularly those requiring complex upper-limb coordination or strong cross-joint coupling. Looking ahead, we will explore richer upper-limb priors, extend to whole-body control, enable online adaptation and personalization, and unify collision and task constraints within a single optimization framework to further improve generalization and robustness in real deployments.